# MTGR: Industrial-Scale Generative Recommendation Framework in Meituan

### v2

@@ Abstract.
Scaling law has been extensively validated in many domains such as natural language processing and computer vision. In the recommendation system, recent work has adopted generative recommendations to achieve scalability, but their generative approaches require abandoning the carefully constructed cross features of traditional recommendation models. We found that this approach significantly degrades model performance, and scaling up cannot compensate for it at all. In this paper, we propose MTGR (Meituan Generative Recommendation) to address this issue. MTGR is modeling based on the HSTU (Zhai et al., 2024) architecture and can retain the original deep learning recommendation model (DLRM) features, including cross features. Additionally, MTGR achieves training and inference acceleration through user-level compression to ensure efficient scaling. We also propose Group-Layer Normalization (GLN) to enhance the performance of encoding within different semantic spaces and the dynamic masking strategy to avoid information leakage. We further optimize the training frameworks, enabling support for our models with 10 to 100 times computational complexity compared to the DLRM, without significant cost increases. MTGR achieved $65$x FLOPs for single-sample forward inference compared to the DLRM model, resulting in the largest gain in nearly two years both offline and online. This breakthrough was successfully deployed on Meituan, the world’s largest food delivery platform, where it has been handling the main traffic.
@@
ccs: Information systems Recommendation systems
@@ 1. Introduction
Scaling law have been proven to apply to most deep learning tasks, including language models(Kaplan et al., 2020), computer vision(Zhai et al., 2022; Peebles and Xie, 2023), and information retrieval(Fang et al., 2024). Under the requirements of high QPS (Queries Per Second) and low latency in industrial recommendation systems, the scaling of the model is usually limited by both the training cost and inference time. Currently, the study on scaling ranking models can be divided into two types: Deep Learning Recommendation Model (DLRM) and Generative Recommendation Model (GRM). DLRM models individual user-item pairs to learn the probability of interest for ranking and scales up by developing more complex mappings. GRM organizes data by token like natural language and performs next token prediction through a transformer architecture.
@@
Unlike the commonly used TensorFlow in the industry, MTGR training framework is built based on TorchRec(Ivchenko et al., 2022) and optimized for computational efficiency. Specifically, to handle the real-time insert/delete of sparse embedding entries, we employ dynamic hash tables to replace static tables. To improve efficiency, we conduct dynamic sequence balancing to address the computation load imbalances among GPUs and adopt embedding ID de-duplication alongside automatic table merging to accelerate embedding lookup. We also incorporate implementation optimization including mixed precision training and operator fusion. Compared to TorchRec, our optimized framework improves training throughput by $1.6$x – $2.4$x while achieving good scalability when running over $100$ GPUs.
We validate the scalability of MTGR on a small-scale dataset. Then, we design three models of varying sizes, which are trained using following over six months of data, to verify the scaling law for both offline and online performance. The large version, compared to the DLRM baseline optimized over years, demonstrates $65$x FLOPs per sample for forward computation and increases the conversion volumes by $1.22\%$ and the CTR (Click-Through Rate) by $1.31\%$ in our scenarios. Meanwhile, the training cost remains unchanged and the inference cost is reduced by 12%. MTGR-large has been deployed in the Meituan take-away recommendation system, serving hundreds of millions of users.
@@ 2.2. Scaling up Recommendation Model
The counterpart to DLRM is GRM. (Zhai et al., 2024) validate the scaling law through HSTU with scaling up to trillion-level parameters. (Deng et al., 2025) use semantic coding to replace traditional ID representations, combining DPO optimization with HSTU to replace the cascaded learning framework with a unified generative model.
@@ 4.1. User Sample Aggregation for Training and Inference Efficiency
Similar, each item $\mathbf{{I}}{}$ in candidates is regarded as a token. The features in candidates are embedded and concatenated, converted to the unified dimension by another MLP. Candidates are concatenated as a series of tokens: $\mathbf{Feat}{}_{\mathbf{{I}}{}}=\text{Concat}(\text{MLP}(\text{Concat}(\mathbf{Emb}{}_{\mathbf{{C}}{}_{i}},\mathbf{Emb}{}_{\mathbf{{I}}{}_{i}})))\in\mathbb{R}^{N_{\mathbf{{I}}{}}}$ Finally, the constructed tokens from ${\mathbf{U}}{}$, $\overrightarrow{\mathbf{S}}{}$, $[\mathbf{{C}}{},\mathbf{s}{}]$ are concatenated, leading a long sequence of tokens:
$\mathbf{Feat}{}_{\mathbb{D}}$ $=$ $\text{Concat}([\mathbf{Feat}{}_{{\mathbf{U}}{}},\mathbf{Feat}{}_{\overrightarrow{\mathbf{S}}{}},\mathbf{Feat}{}_{\overrightarrow{\mathbf{R}}{}},\mathbf{Feat}{}_{\mathbf{{I}}{}}])$
@@ 4.2. Unified HSTU Encoder
Similar in LLM, the input token sequences are processed layer-wisely. As shown in Fig.2, in the self-attention block, input sequence $\mathbf{X}$ is firstly normalized by group layer norm. Features of the same domain (for example, ${\mathbf{U}}{}$) form a group. Then, the group layer norm ensures that tokens from different domains share a similar distribution before self-attention and aligns different semantic spaces of different domains $\tilde{X}=\text{GroupLN}(\mathbf{X})$. The normalized input is then projected to 4 different representation: $\mathbf{K},\mathbf{Q},\mathbf{V},\mathbf{U}=\text{MLP}_{\mathbf{K}/\mathbf{Q}/\mathbf{V}/\mathbf{U}}(\tilde{\mathbf{X}})$. $\mathbf{Q}$, $\mathbf{K}$ is utilized for multi-head attention computation with silu non-linear activation. The leading attention is divided by total length of input features as an average factor. Next, a customized mask ($\mathbf{M}$) is imposed on the attention score and the projected $\mathbf{V}$ is applied for value update:
@@ 5. Training System
Dynamic Hash Table. TorchRec use a fixed-size table to process sparse embeddings, which is not suitable for large-scale industrial streaming training scenarios. Firstly, additional embeddings cannot be allocated to new users and items at real-time once the static table hits its preset capacity. Although some default embeddings may be used in this case, or some eviction strategies can be used to eliminate ‘old’ IDs, model accuracy will be degraded. Secondly, static tables typically require pre-allocation of capacity exceeding actual requirements to prevent ID overflow, inevitably leading to inefficient memory utilization. To tackle the challenges, we implement a hash-based and high performance embedding table, dynamically allocating space for sparse IDs during the training process. We employ a decoupled architecture for hash table, which separates key and value storage into distinct structures. The key structure maintains a lightweight mapping table containing keys and corresponding pointers to embedding vectors, while the value structure stores both the embedding vectors and auxiliary metadata (e.g., counters and timestamps) required for eviction policies. This dual-structure design achieves two critical objectives: (1) enabling dynamic capacity expansion through replication of only the compact key structure rather than huge embedding data, and (2) optimizing key scanning efficiency by storing keys in compact memory layouts while accommodating potentially sparse key distributions.
Embedding Lookup. The embedding lookup process employs All-to-all communication for cross-device embedding exchange. We adopt two-stage ID unique operation to reduce redundant IDs before and after ID communication to avoid repetitive embedding transfers among the devices. We also design a feature configuration interface to implement automatic table merging, which can reduce the number of embedding lookup operators and thus speed up the process.
@@
Other Optimizations. To further enhance training efficiency, we maximize parallelism through pipeline technology using three streams: copy, dispatch, and compute. Specifically, the copy stream loads input data from CPU to GPU, the dispatch stream performs table lookups based on IDs, and the compute stream handles forward computation and backward updates. For instance, while the compute stream executes forward and backward passes for batch $\mathbf{T}$, the copy stream concurrently loads batch $\mathbf{T}+1$ to mask I/O latency. Upon completing backward updates for batch $\mathbf{T}$, the dispatch stream immediately initiates table lookups and communication for batch $\mathbf{T}+1$. Moreover, we employ bf16 mixed-precision training and develop a well-crafted cutlass-based attention kernel which is similar to FlashAttention to speed up the training process.
@@ 6.1. Experiment setup
Datasets. Public datasets widely use independent ID and attribute features, where cross features are seldom introduced. However, cross features demonstrate importance in real-world application. Cross features are an important category of features in our scenario. They are often meticulously hand-crafted, including interactions like user-item, user and higher-level categories, item and spatio-temporal information. To compensate the absence of cross feature in public datasets, we construct a training dataset based on logs from the real industrial-scale recommendation system in Meituan. Unlike public datasets, our real dataset contains a richer cross features set and longer user behavior sequences. In addition, the volume of our dataset is large, allowing complex models to achieve more adequate convergence during training. For the offline experiments, we collect data over a $10$-day period. The statistics of the dataset are shown in Table  1. For the online experiments, in order to compare with the DLRM baseline that has been trained for over $2$ years, we constructed a longer-term dataset for the experiments, using data spanning more than $6$ months.
@@
Parameter Setting. Our model is trained using the Adam optimizer. For DLRM, each GPU processes a batch size of 2400, with 8 NVIDIA A100 GPUs for training. In the case of MTGR, the batch size is set to 96, utilizing 16 NVIDIA A100 GPUs for training. As shown in Table 2, the learning rate decreases as the complexity of the model increases. Moreover, as the model size grows, we proportionally increase the size of the sparse parameters. Assuming that a token consists of $k$ features, the embedding dimension for each feature is generally set to an integer close to $d_{\text{model}{}}/k$. Finally, the maximum length of $\overrightarrow{\mathbf{S}{}}$ is set to $1000$, while $\overrightarrow{\mathbf{R}}{}$ is set to $100$.
@@ 6.2. Overall Performance Comparison
Table 3. Overall performance. Impr.represents the relative improvement of the best MTGR model compared to the strongest DLRM baselines (underlined).

### v4

@@ Abstract.
Scaling law has been extensively validated in many domains such as natural language processing and computer vision. In the recommendation system, recent work has adopted generative recommendations to achieve scalability, but their generative approaches require abandoning the carefully constructed cross features of traditional recommendation models. We found that this approach significantly degrades model performance, and scaling up cannot compensate for it at all. In this paper, we propose MTGR (Meituan Generative Recommendation) to address this issue. MTGR is modeling based on the HSTU (Zhai et al., 2024) architecture and can retain the original deep learning recommendation model (DLRM) features, including cross features. Additionally, MTGR achieves training and inference acceleration through user-level compression to ensure efficient scaling. We also propose Group-Layer Normalization (GLN) to enhance the performance of encoding within different semantic spaces and the dynamic masking strategy to avoid information leakage. We further optimize the training frameworks, enabling support for our models with 10 to 100 times computational complexity compared to the DLRM, without significant cost increases. MTGR achieved 6565x FLOPs for single-sample forward inference compared to the DLRM model, resulting in the largest gain in nearly two years both offline and online. This breakthrough was successfully deployed on Meituan, the world’s largest food delivery platform, where it has been handling the main traffic.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761565isbn: 979-8-4007-2040-6/2025/11ccs: Information systems Recommendation systems
@@ 1. Introduction
Scaling law have been proven to apply to most deep learning tasks, including language models(Kaplan et al., 2020), computer vision(Zhai et al., 2022; Peebles and Xie, 2023), and information retrieval(Fang et al., 2024). Our work focuses on scaling up the ranking model in industrial recommendation systems effectively. Under the requirements of high QPS (Queries Per Second) and low latency in industrial recommendation systems, the scaling of the model is usually limited by both the training cost and inference time. Currently, the study on scaling ranking models can be divided into two types: Deep Learning Recommendation Model (DLRM) and Generative Recommendation Model (GRM). DLRM models individual user-item pairs to learn the probability of interest for ranking and scales up by developing more complex mappings. GRM organizes data by token like natural language and performs next token prediction through a transformer architecture.
@@
Unlike the commonly used TensorFlow in the industry, MTGR training framework is built based on TorchRec(Ivchenko et al., 2022) and optimized for computational efficiency. Specifically, to handle the real-time insert/delete of sparse embedding entries, we employ dynamic hash tables to replace static tables. To improve efficiency, we conduct dynamic sequence balancing to address the computation load imbalances among GPUs and adopt embedding ID de-duplication alongside automatic table merging to accelerate embedding lookup. We also incorporate implementation optimization including mixed precision training and operator fusion. Compared to TorchRec, our optimized framework improves training throughput by 1.61.6x – 2.42.4x while achieving good scalability when running over $100$ GPUs.
We validate the scalability of MTGR on a small-scale dataset. Then, we design three models of varying sizes, which are trained using following over six months of data, to verify the scaling law for both offline and online performance. The large version, compared to the DLRM baseline optimized over years, demonstrates 6565x FLOPs per sample for forward computation and increases the conversion volumes by $1.22\%$ and the CTR (Click-Through Rate) by $1.31\%$ in our scenarios. Meanwhile, the training cost remains unchanged and the inference cost is reduced by 12%. MTGR-large has been deployed in the Meituan take-away recommendation system, serving hundreds of millions of users.
@@ 2.2. Scaling up Recommendation Model
The counterpart to DLRM is GRM. (Zhai et al., 2024) validate the scaling law through HSTU with scaling up to trillion-level parameters. (Deng et al., 2025) use semantic coding to replace traditional ID representations, combining DPO optimization with a transformer-based framework to replace the cascaded learning framework with a unified generative model.
@@ 4.1. User Sample Aggregation for Training and Inference Efficiency
Similar, each item $\mathbf{{I}}{}$ in candidates is regarded as a token. The features in candidates are embedded and concatenated, converted to the unified dimension by another MLP. Candidates are concatenated as a series of tokens: $\mathbf{Feat}{}_{\mathbf{{I}}{}}=\text{Concat}(\text{MLP}(\text{Concat}(\mathbf{Emb}{}_{\mathbf{{C}}{}_{i}},\mathbf{Emb}{}_{\mathbf{{I}}{}_{i}})))\in\mathbb{R}^{N_{\mathbf{{I}}{}}}$ Finally, the constructed tokens from $\mathbf{U}$}{}, $\overrightarrow{\mathbf{S}}{}$, $\mathbf{{C}}{},\mathbf{s}{}]$ are concatenated, leading a long sequence of tokens:
[…]
@@ 4.2. Unified HSTU Encoder
[…]
@@ 5. Training System
Dynamic Hash Table. TorchRec use a fixed-size table to process sparse embeddings, which is not suitable for large-scale industrial streaming training scenarios. Firstly, additional embeddings cannot be allocated to new users and items at real-time once the static table hits its preset capacity. Secondly, static embedding tables often require reserving more space than needed to avoid ID overflow, resulting in ineffective use of memory resources. To address these issues, we develop a high-performance embedding table utilizing hash techniques, enabling the dynamic allocation of space to accommodate sparse IDs during training. Our design employs a decoupled hash table architecture(Wang et al., 2025), segregating key storage and value storage into separate entities. The key storage offers a lightweight mapping system linking keys to pointers, which point to embedding vectors, while the value storage contains the embedding vectors along with additional metadata, such as counters and timestamps, for eviction policies. This two-part system achieves two primary objectives: (1) facilitating dynamic expansion of capacity by replicating only the key storage instead of the extensive embeddings, and (2) improving the efficiency of key scanning by arranging keys in a compact format.
Embedding Lookup. The embedding lookup process employs All-to-all communication for cross-device embedding exchange. To minimize duplicate ID transfers between devices, we implement a two-step process that ensures IDs remain unique both before and after communication.
@@
Other Optimizations. To further enhance training efficiency, we implement pipeline technology utilizing three distinct streams: copy, dispatch, and compute. The copy stream is responsible for transferring input data from the CPU to the GPU, while the dispatch stream executes table lookups using IDs, and the compute stream deals with both forward computations and backward updates. For example, while the compute stream performs forward and backward passes for batch $\mathbf{T}$, the copy stream simultaneously loads batch $\mathbf{T}+1$, thereby minimizing I/O delay. Once backward updates for batch $\mathbf{T}$ are completed, the dispatch stream promptly starts table lookups and communication for batch $\mathbf{T}+1$. Furthermore, we employ bf16 mixed-precision training and have designed a specialized attention kernel based on cutlass to accelerate training progress.
@@ 6.1. Experiment setup
Datasets. Public datasets widely use independent ID and attribute features, where cross features are seldom introduced. However, cross features demonstrate importance in real-world application. Cross features are an important category of features in our scenario. They are often meticulously hand-crafted, including interactions like user-item, user and higher-level categories, item and spatio-temporal information. To compensate the absence of cross feature in public datasets, we construct a training dataset based on logs from the real industrial-scale recommendation system in Meituan. Unlike public datasets, our real dataset contains a richer cross features set and longer user behavior sequences. Using our industrial-scale data for experiments better highlights the significant impact of these cross features on real recommendation systems. In addition, the volume of our dataset is large, allowing complex models to achieve more adequate convergence during training. For the offline experiments, we collect data over a $10$-day period. The statistics of the dataset are shown in Table  1. For the online experiments, in order to compare with the DLRM baseline that has been trained for over $2$ years, we constructed a longer-term dataset for the experiments, using data spanning more than $6$ months.
@@
Parameter Setting. Our model is trained using the Adam optimizer. For DLRM, each GPU processes a batch size of 2400, with 8 NVIDIA A100 GPUs for training. In the case of MTGR, the batch size is set to 96, utilizing 16 NVIDIA A100 GPUs for training. As shown in Table 2, the learning rate decreases as the complexity of the model increases. Additionally, as the computational complexity grows, we proportionally enlarge the size of the sparse parameters by configuring different embedding dimensions. Assuming that a token consists of $k$ features, the embedding dimension for each feature is generally set to an integer close to $d_{\text{model}{}}/k$. It is worth noting that, to prevent excessive overhead due to the overexpansion of sparse parameters, we primarily increase the dimensions of sparse features with smaller cardinality, while maintaining the dimensions of extremely sparse features unchanged. Finally, the maximum length of $\overrightarrow{\mathbf{S}{}}$ is set to $1000$, while $\overrightarrow{\mathbf{R}}{}$ is set to $100$.
@@ 6.2. Overall Performance Comparison
Table 3. Overall performance. Impr.% represents the relative improvement of the best MTGR model compared to the strongest DLRM baselines (underlined).